ON K_{∗}(ℤ) AND CLASSICAL CONJECTURES IN

THE ARITHMETIC OF CYCLOTOMIC FIELDS

by Wojciech Gajda

Abstract. This paper contains a short survey on the problem of computing Kgroups of the integers and its relations to the classical conjectures in number theory, such as the Vandiver conjecture. The main point which we want to stress is that the topological methods of K-theory can be used in proving new theorems in the direction of the number theoretical conjectures.

0. Introduction

The algebraic K-groups of schemes introduced by Quillen are expected to carry deep arithmetic information, according to conjectures of Lichtenbaum, Beilinson, Bloch and Kato which relate K-theory to special values of L-functions. Despite the substantial progress which has been achieved, there are still many features of the relationship of the K-theory with number theory which remain to be understood.

One of the very fascinanting aspects of algebraic K-theory is the link between Kgroups of ℤ and deep questions in number theory, such as the Vandiver conjecture. The conjecture asserts that for any prime number l, the class number h_{l}^{+} of the maximal totally real subfield ℚ(cos \frac{2π}{l}) of the field of lth roots of unity, is prime to l. The statement of the conjecture is one of the most important questions in the theory of cyclotomic fields.

It comes as a surprise that the Vandiver conjecture - a classical problem in number theory, can be reformulated in terms of K-groups, whose definition and properties come from algebraic topology. Even more surprisingly, it turns out that using K-theory of ℤ one can prove new results in the direction of the conjecture. One of our goals in this paper is to present some of the recently proven results of this type, due to Banaszak, Kurihara, Soulé and the author (cf. Theorems: 4.4, 4.7, 4.8 and 5.5 below).

After recalling some basic facts on the K-theory of number fields, in Section 2 we review the current state of the knowledge on the K-groups of the integers,

Typeset by AMS-TEX

1991 Mathematics Subject Classification. Primary 19F27; Secondary 19Fxx.

the conjecture of Kurihara on K_{∗}(ℤ) and relations to the conjectures of Quillen-Lichtenbaum and Bloch-Kato. In Section 3 we introduce the Beilinson element which is a nontorsion element of K_{2n+1}(ℤ), where n is an even natural number. The Vandiver conjecture can be reformulated in terms of the Beilinson element. The conjecture is true if and only if for all n as above, the group K_{2n+1}(ℤ) modulo 2-torsion, is generated by the Beilinson element. In Sections 4 and 5, in addition to the theorems mentioned above, we discuss other reformulations of the Vandiver conjecture and of the weaker form, due to Iwasawa.

In the moment one of the most important problems concerning K-groups of the integers is to construct nonzero elements (generators) in the groups K_{2n}(ℤ) for n odd (cf. Remark 5.6). One can expect that such a construction can be provided by a topological method.

Acknowledgements: The paper contains an expanded version of my lectures on the subject given at Northwestern University during the Emphasis Year in Algebraic Topology in 2002. It is my pleasure to thank the organizers of the Year: E.Friedlander, P.Goerss and S.Priddy for inviting me to participate in the activity. The paper was written up while I was visiting the Max-Planck-Institute für Mathematik in Bonn. I would like to acknowledge the financial support of the MPI.

In Poznań, the author was partially supported by a KBN grant 2 P03A 048 22.

1. Basic facts on K-groups of number fields

In the paper [Q1] Quillen introduced higher K-groups of a commutative ring R with unity as the homotopy groups:

K_{m}(R) = π_{m+1}(BQP(R))

for any m ≥ 0 cf. [Q1, Def., p.103 and Example, p.104]. Here P(R) denotes the category of finitely generated, projective R-modules, Q(−) stands for the categorical Q-construction and BC is the classifying space of a small category C. Substituting m=0, 1, 2 into the definition we recover the groups: K_{0}(R), K_{1}(R) and K_{2}(R) which were introduced earlier by Grothendieck, Bass and Milnor, respectively. Recall that

K_{0}(R) = F (R)/N(R),

where F (R) is the free abelian group generated by the isomorphism classes of Rmodules from P(R) and N(R) denotes the subgroup generated by [P ⊕Q]−[P]−[Q], for P, Q ∈ P(R). The group K_{1} of Bass is by definition:

K_{1}(R) = GL(R)/E(R) = H_{1}(GL(R), ℤ)

ON K_{∗}(ℤ) AND CLASSICAL CONJECTURES 3

the abelianization of the general linear group GL(R)= ⋃_{n≥0} GL_{n}(R), where we identify Gl_{n}(R) with its image in GL_{n+1}(R) by sending an invertible n × n matrix A to ( \begin{matrix}A & 0 \\ 0 & 1 \end{matrix} ). The abelianization of GL coincides with the first homology group,

because for any discrete group G we have π_{1}(BG)=G. The group K_{2}(R) of Milnor (usually defined in terms of Steinberg symbols) is the second homology H_{2}(E(R), ℤ) of the group of elementary matrices. Proofs of the following facts can be found in the textbooks on K-theory cf: [M] and [R].

Theorem 1.1.

(1) If L is a field, then K_{0}(L)=ℤ and K_{1}(L)=L^{×} (the multiplicative group of L). Moreover we have (Matsumoto’s theorem):

K_{2}(L) = (L^{×} ⊗_{ℤ} L^{×})/N,

where

N = ⟨x ⊗ (1−x) : x ∈ L, x ≠ 0, 1⟩

is the subgroup generated by the Steinberg relations.

(2) For the ring of integers O_{F} of the number field F:

K_{0}(O_{F})=Cl(O_{F}) ⊕ ℤ  and  K_{1}(O_{F})=O_{F}^{×},

where Cl(O_{F}) is the ideal class group of the ring O_{F} and O_{F}^{×} denotes its group of invertible elements (the group of global units of O_{F}). The group K_{2}(O_{F}) is finite and can be explicitly computed in some special cases.

The class group and the group of units are two of the most important arithmetical invariants of the ring O_{F} . The class group can be defined by the exact sequence:

(1.2)  1 ⟶ O_{F}^{×} ⟶ F^{×} ⟶^{∂} ⨁_{𝔭} ℤ ⟶ Cl(O_{F}) ⟶ 0

where the summation is taken over all prime ideals of O_{F} . The map ∂ = ⨁ ∂_{𝔭} is defined by the valuations. For c ∈ O_{F} we define ν_{𝔭}(c), to be the maximal integer such that the ideal 𝔭^{ν_{𝔭}(c)} divides the principal ideal cO_{F} . We put:

∂_{𝔭}(\frac{a}{b}) = ν_{𝔭}(a)−ν_{𝔭}(b)

where a, b ∈ O_{F} , and b ≠ 0 (this clearly makes sense, because F is the field of fractions of the ring O_{F}). Using Theorem 1.1 the exact sequence (1.2) can be identified with: (1.3)

0 ⟶ K_{1}(O_{F}) ⟶ K_{1}(F) ⟶^{∂} ⨁_{𝔭} K_{0}(κ_{𝔭}) ⟶ K_{0}(O_{F})/ℤ ⟶ 0

where κ_{𝔭}=O_{F} /𝔭 is the residue field at 𝔭. The residue field is finite. It follows by [Q1, Cor., p.113] that the sequence (1.3) extends to the exact localization sequence:

⟶ K_{m}(O_{F}) ⟶ K_{m}(F) ⟶^{∂} ⨁_{𝔭} K_{m−1}(κ_{𝔭}) ⟶ K_{m−1}(O_{F}) ⟶

for any m ≥ 1. Using properties of the map ∂ and the vanishing of even K-groups of finite fields (proven by Quillen in [Q2]) Soulé showed that the localization sequence breaks down into short exact sequences.

Theorem 1.4. Let n > 0.

(1) There exists a short exact sequence

0 ⟶ K_{2n}(O_{F}) ⟶ K_{2n}(F) ⟶^{∂} ⨁_{𝔭} K_{2n−1}(κ_{𝔭}) ⟶ 0.

(2) There is an isomorphism K_{2n+1}(F) ≅ K_{2n+1}(O_{F}).

Another important result of Quillen concerns the K-groups of algebraic integers.

Theorem 1.5. [Q3] For any number field F the groups K_{m}(O_{F}) are finitely generated.

Borel computed ranks of these groups using cohomology of arithmetic groups.

Theorem 1.6. [Bo]

rank_{ℤ} K_{m}(O_{F}) = \{ \begin{matrix}1 & m = 0 \\ r_{1} + r_{2} − 1 & m = 1 \\ 0 & m > 0, even \\ r_{1} + r_{2} & m > 1, and m ≡ 1 mod 4 \\ r_{2} & m ≡ 3 mod 4, \end{matrix}

where r_{1} (2r_{2}, resp.) denotes the number of real (complex, resp.) embeddings of the field F, hence r_{1}+2r_{2} = [F : ℚ].

Note that the rank of the group K_{2n+1}(O_{F}) equals the order of vanishing of the Dedekind zeta function of F at s = −n. The celebrated Lichtenbaum conjecture [Li] postulates formulas for values of the Dedekind zeta at odd negative integers in terms of the K-groups of O_{F} for a totally real number field F (so then r_{2}=0). In the case of F =ℚ (which is of the main interest for us here) O_{F} =ℤ, r_{1}=1, r_{2}=0 and according to the Lichtenbaum conjecture we should have:

(1.7)  |ζ(−n)| = 2\frac{#K_{2n}(ℤ)}{#K_{2n+1}(ℤ)},

for any odd natural number n, where ζ(s) is the Riemann zeta function. This formula follows from the Quillen-Lichtenbaum conjecture (cf. section 3) as was shown by Kolster et al. in [KFN1] and [KFN2]. The Quillen-Lichtenbaum conjecture follows if the Bloch-Kato conjecture is true (cf. Remark 2.2 for the discussion).

Dealing with torsion in K-groups, it is very useful to have K-theory with finite coefficients introdued by Karoubi and Browder cf. [Br]. For natural numbers n and m ≥ 1 we define

K_{m}(R, ℤ/n) = π_{m+1}(BQP(R), ℤ/n)

as the homotopy group with coefficients in ℤ/n cf. [N]. The homotopy group π_{m}(X, ℤ/n) of a topological space is the set of homotopy classes of maps from the Moore space M^{m}(ℤ/n) to X. The Moore space M^{m}(ℤ/n) is the mapping cone of the map S^{m−1} → S^{m−1} which induces mutliplication by n on π_{m}. We will use K-groups with coefficients in ℤ/l^{k}, for an odd prime l and k ≥ 1. The cofibration sequence in homotopy induces short exact sequences:

0 ⟶ K_{m}(R)/l^{k} ⟶ K_{m}(R, ℤ/l^{k}) ⟶^{B} K_{m−1}(R)[l^{k}] ⟶ 0

where we denote by C[l^{k}] the subgroup of l^{k}-torsion elements of an abelian group C. The map B is called the Bockstein map. Clearly, the K-groups with ℤ/l^{k} are torsion groups and are anihilated by l^{2k}. Let R be a Dedekind domain, which contains a primitive root of unity ξ_{l^{k}} of order l^{k}, e.g., let R = ℤ[ξ_{l^{k}}]. Consider the surjective Bockstein map:

(1.8)  B: K_{2}(R, ℤ/l^{k}) → K_{1}(R)[l^{k}]

Observe that: K_{1}(R)[l^{k}] = R^{×}[l^{k}]. Let μ_{l^{k}} denote the subgroup of l^{k}th roots of unity in R^{×}. There exists the Bott element β = β(ξ_{l^{k}}) ∈ K_{2}(R, ℤ/l^{k}) (which depends only on ξ_{l^{k}}), such that B(β(ξ_{l^{k}})) = ξ_{l^{k}}. By definition, β(ξ_{l^{k}}) is the image of ξ_{l^{k}} by the composition of maps:

μ_{l^{k}} ⟶^{≅} π_{1}(BGL_{1}(R))[l^{k}] ⟶^{≅} π_{2}(BGL_{1}(R), ℤ/l^{k}) → K_{2}(R, ℤ/l^{k}).

In K-theory and K-theory with finite coefficients there exist transfer maps similar to the transfer maps in group cohomology cf. [Q1, p. 111] and [So1]. In the case of number fields we have the localization sequence in K-theory with finite coefficients:

\begin{matrix}⟶ & K_{m}(O_{F} , ℤ/l^{k}) & ⟶ & K_{m}(F, ℤ/l^{k}) & ⟶^{∂} \\ ⟶ & ⨁_{𝔭 |̸l} K_{m−1}(κ_{𝔭}, ℤ/l^{k}) & ⟶ & K_{m−1}(O_{F} , ℤ/l^{k}) & ⟶ \end{matrix}

which doesn’t split, in general. There are products maps:

⋆ : K_{i}(R) ⊗ K_{j}(R) → K_{i+j}(R)

which were constructed by Loday [Lo] (see also [We1]). When i=j=0, then the product is induced by the tensor product of R-modules. Product maps in K-theory are constructed using multiplicative properties of the space BQP(R), which is the zero space of an E_{∞}-spectrum in the sense of algebraic topology cf. [Wal]. All expected properties of the products hold, which makes ⨁_{i≥0} K_{i}(R) into a graded commutative ring. For K-groups with finite coefficients, the product maps:

⋆ : K_{i}(R, ℤ/l^{k}) ⊗ K_{j}(R, ℤ/l^{k}) → K_{i+j}(R, ℤ/l^{k}),

(at least for l odd) were constructed by Browder in [Br]. Note that there are no well behaved product maps for l = 2 and k ≤ 3 cf. loc. cit.

2. What we know about K_{∗}(ℤ) today

If F = ℚ and m ≥ 2, then Borel’s theorem implies:

rank K_{m}(ℤ) = \{ \begin{matrix}1 & if m ≡ 1 mod 4 \\ 0 & otherwise \end{matrix}

With exception for K_{0}(ℤ) = ℤ and K_{1}(ℤ) = ℤ/2, the computation of the K-groups of the integers turned out to be quite a difficult problem. As for today the following K-groups of ℤ are known:

(a) K_{2}(ℤ) = ℤ/2 (see [M])

(b) K_{3}(ℤ) = ℤ/48 (Lee and Szczarba, [LS1] )

(c) K_{4}(ℤ) = 0 (Lee, Szczarba and Soulé considered the odd torsion cf. [LS2] and [So6]. Rognes showed that the group has no 2-torsion cf. [R])

(d) K_{5}(ℤ) = ℤ (cf. [LS2] and [E-VGS])

(e) K_{m}(ℤ)⊗ ℤ_{2}, for m ≥ 6, is known due to computations of Rognes and Weibel cf. [RW] and [We2], who used the Milnor conjecture proven by Voevodsky, [Vo1]. By ℤ_{l} we denote here the ring of l-adic integers, for a prime number l.

There are precise conjectures concerning torsion of K_{m}(ℤ) which were formulated by Kurihara cf. [K] and also, independently by Mitchell.

Kurihara Conjecture. Up to 2-torsion, the following statements are true:

(1) K_{2n}(ℤ) = ℤ/N_{n} for n > 0, odd

(2) K_{2n+1}(ℤ) = ℤ/D_{n} for n > 0, odd

(3) K_{2n}(ℤ) = 0 for n > 0, even

(4) K_{2n+1}(ℤ) = ℤ for n > 0, even.

Here N_{n} and D_{n} are such relatively prime natural numbers that

|ζ(−n)| = \frac{N_{n}}{D_{n}} ,

where ζ(s) is the Riemann zeta function. Euler showed that for positive, odd integers n

ζ(−n) = −\frac{B_{n+1}}{n+1} ,

where B_{k} denote the Bernoulli numbers, which are defined by the generating series

\frac{t}{e^{t} − 1} = ∑_{j≥0} \frac{B_{j}}{j!} t^{j}

and can be computed by the recursive formula:

\frac{B_{m}}{m!} = − ∑_{j=0}^{m−1} \frac{B_{j}}{j!} \frac{1}{(m−j+1)!}.

It is not difficult to check that B_{0}=1, B_{1}= − \frac{1}{2} and B_{j}=0, for odd j > 1. Moreover one computes (see the tables in [Wa]):

B_{2} = \frac{1}{6}, B_{4} = −\frac{1}{30}, B_{6} = \frac{1}{42}, B_{8} = −\frac{1}{30}, B_{10} = \frac{5}{66}, B_{12} = −\frac{691}{2730}
B_{14} = \frac{7}{6}, B_{16} = −\frac{3617}{510}, B_{18} = \frac{43867}{798}, B_{20} = −\frac{174611}{330}, B_{22} = \frac{854513}{138}.

Thus the first numbers D_{n} are:

D_{1} = 12, D_{3} = 120, D_{5} = 252, D_{7} = 240, D_{9} = 132, D_{11} = 32760,
D_{13} = 12, D_{15} = 8160, D_{17} = 14364, D_{19} = 6600, D_{21} = 276.

The denominators D_{n} can be computed completely due to the classical theorem of von Staudt [Wa, App.B, Th. B4], which says that a prime power p^{a} divides D_{n}, if and only if, (p−1)p^{a−1} divides n+1. The numerators N_{n} are much more mysterious. For example, according to (1) of the Kurihara’s conjecture K_{38}(ℤ)=ℤ/174611 up to 2-torsion. Note that the statements (1) and (2) of the conjecture are consistent with the Lichtenbaum conjecture for ℚ. According to (3) and (4) the groups:

K_{8}(ℤ), K_{12}(ℤ), K_{16}(ℤ), K_{20}(ℤ), … should vanish,
K_{9}(ℤ), K_{13}(ℤ), K_{17}(ℤ), K_{21}(ℤ), … should be isomorphic to ℤ,

if we ignore the 2-torsion.

K-groups of the integers map to Galois cohomology by Chern character maps.

Theorem 2.1. [Soulé, Dwyer and Friedlander] For any odd prime l and any n > 0, there exist epimorphisms of groups

(1) K_{2n}(ℤ) ⊗ ℤ_{l} ⟶ H^{2}(G, ℤ_{l}(n+1))

(2) K_{2n+1}(ℤ) ⊗ ℤ_{l} ⟶ H^{1}(G, ℤ_{l}(n+1)).

Here G=Gal(ℚ_{S}/ℚ) is the Galois group (with its natural profinite topology) of the maximal field extension ℚ_{S}/ℚ contained in ℚ̄ and such that ℚ_{S} is unramified outside of the principal ideal (l). The group acts on all the l-primary roots of unity, since all the cyclotomic fields ℚ(ξ_{l^{k}}) are contained in ℚ_{S}. The coefficient module ℤ_{l}(n+1) is the (n+1)th tensor power of the G-module

ℤ_{l}(1)= lim_{←} ℤ/l^{k}(1),

which is the inverse limit of the G-modules of l^{k}th roots of unity. The cohomology is defined by continuous cochains using the topology of G and the compact Gmodule ℤ_{l}(n+1). The Dwyer-Friedlander maps were defined by methods of algebraic topology see [DF], Prop. 4.4. Surjectivity follows by [DF], Th. 8.7 and Rem. 8.8. Note that the maps constructed in [DF] take values in groups of continuous étale cohomology H^{i}(ℤ[\frac{1}{l}]; ℤ_{l}(n + 1)). We have replaced the étale cohomology by the isomorphic H^{i}(G, ℤ_{l}(n+1)), (cf. [Mi, p. 209] or [Sch, p. 204]).

Quillen-Lichtenbaum Conjecture. The maps from Theorem 2.1 are isomorphisms.

Remark 2.2. This is a special case of the conjecture which was formulated for the ring of integers of any number field. It was proven by Levine cf. [Le1] and [Le2] (see also [SV] and [GL]) that the Quillen-Lichtenbaum conjecture follows from the Bloch-Kato conjecture, which says that the norm residue map

K_{m}^{M}(L)/l ⟶ H^{m}(G_{L}, ℤ/l(m))

from the Milnor K-theory to Galois cohomology is an isomorphism, for any field L of characteristic prime to l. The Milnor K-group K_{m}^{M}(L) is defined as the quotient of the mth tensor power (L^{×})^{⊗m} by the submodule generated by Steinberg relations, similar to the subgroup N of Theorem 1.1 (1) for m=2. The norm residue map for m=1 is the isomorphism:

L^{×}/L^{×l} ⟶^{≅} H^{1}(G_{L}, ℤ/l(1)).

It is obtained when one takes cohomology of the exact sequence of G_{L}-modules:

0 ⟶ ℤ/l(1) ⟶ L̄^{×} ⟶^{×l} L̄^{×} ⟶ 1,

because H^{0}(G_{L}, L̄^{×})=L^{×} and H^{1}(G_{L}, L̄^{×}) vanishes by the Hilbert Theorem 90 cf. [La, Th. 10.1]. For m>1 the norm residue map is defined by sending an element {a_{1}, a_{2}, …, a_{m}} of K_{m}^{M}(L) to the cup product a_{1} ∪ a_{2} ∪ ⋯ ∪ a_{m} in Galois cohomology. There are strong reasons to believe that the proof of the Bloch-Kato conjecture for odd primes will be available soon. Then the Quillen-Lichtenbaum and the Lichtenbaum conjecture (at least for ℤ) will be proven, too. The case of l = 2 of the Bloch-Kato conjecture, i.e., the Milnor conjecture, was proven by Voevodsky in [Vo1].

Remark 2.3. The Galois cohomology groups which appear in Theorem 2.1 belong more to number theory than to topology. The current knowledge of these groups, especially of H^{2}:=H^{2}(G, ℤ_{l}(n+1)), is limited.

• The group H^{2} is finite. It is an immediate corollary of the finiteness of K_{2n}(ℤ) and the surjectivity of the Dwyer-Friedlander map.

• If n is even and positive the group H^{2} should vanish by (3) of the conjecture of Kurihara and Theorem 2.1. The vanishing is consistent with the Vandiver conjecture cf. section 4.

• If n is odd, it follows by the Main Conjecture in Iwasawa theory proven for ℚ by Mazur and Wiles in [MW], that the group H^{2} has l^{k} elements, where l^{k} is the l-part of the number N_{n}. Theorem 2.1 and part (1) of the Kurihara conjecture imply that H^{2} should be cyclic. The question about the cyclicity of H^{2}, for n odd, is related to another conjecture in number theory, which we will discuss in section 5.

• For odd n, the group H^{1}:=H^{1}(G, ℤ_{l}(n+1)) has l^{k} elements and is cyclic, where l^{k} is the l-part of D_{n}. It is easy to prove this fact using standard Galois cohomology. In this case the Quillen-Lichtenbaum conjecture and the results on the image of J-homomorphism (see [Q4] and [Br, Th. 4.8]) give explicit generators of the group K_{2n+1}(ℤ).

Remark 2.4. On the side of K_{∗}(ℤ), if the Dwyer-Friedlander maps are proven to be isomorphisms, then there is still plenty to do.

• As mentioned above, the part (2) of Kurihara’s conjecture follow if the Quillen-Lichtenbaum conjecture is proven. To prove part (1) we will still need a generator in the group K_{2n}(ℤ), for n odd.

• Not much is known in the moment about (3) and (4). Note that (3) and (4) are equivalent cf. Proposition 4.11, if the Quillen-Lichtenbaum conjecture holds true. Any results in the direction of statements: (1), (3) or (4) of the Kurihara conjecture will be of upmost interest for number theory.

Quillen in the classical paper [Q2] computed K-groups of finite fields using methods of algebraic topology.

Theorem 2.5.

K_{m}(𝔽_{q}) = \{ \begin{matrix}0 & if m > 0 and m is even \\ ℤ/(q^{n+1}−1) & if m = 2n+1, \end{matrix}

where 𝔽_{q} denotes the field with q elements, and q is a power of a prime number.

3. Special elements in K_{2n+1}(ℤ)

The reader has seen already that it is quite difficult to compute K-groups of the integers. Using Quillen’s calculation of K-groups of finite fields one can try to compare K_{∗}(ℤ) with the groups K_{∗}(𝔽_{p}). In this section we discuss such a comparision given by the map

K_{2n+1}(ℤ) ⟶ K_{2n+1}(𝔽_{p})

induced by reductions at prime numbers p, for n even, positive. In this case the conjecture of Kurihara predicts that K_{2n+1}(ℤ) modulo 2-torsion is ℤ. Sketching the proof of Theorem 3.2 below we introduce into discussion the Beilinson element which conjecturally generates the group K_{2n+1}(ℤ). We also introduce certain cyclotomic numbers, which will be particularly useful in the next section when we talk about the Vandiver conjecture.

Definition 3.1. Let P be a set of rational prime numbers. We say that P has natural density d(P), if the limit

lim_{x→∞} \frac{#\{p ≤ x: p ∈ P\}}{#\{p ≤ x\}}

exists and equals d(P). It is clear from the definition that if P has density d, then 0 ≤ d ≤ 1 and that every finite set of primes has density 0.

Theorem 3.2. Let n be an even, positive integer. Assume that l is an odd prime which does not divide the number (n+1)b, where b is an integer (depending on n only) defined in the proof. Let k be a fixed positive integer. Denote by M_{k} the set of rational primes p which satisfy the following two conditions:

(1) the l-torsion part of the group K_{2n+1}(𝔽_{p}) is isomorphic to ℤ/l^{k}

(2) the reduction map K_{2n+1}(ℤ) ⟶ K_{2n+1}(𝔽_{p})_{l} is nontrivial.

Then the set M_{k} has positive density, which equals \frac{l^{k}−1}{l^{2k}}.

Proof. Step 1. Cyclotomic elements of Soulé and the element of Beilinson Recall that, since the K-groups of the integers are finitely generated, we have:

K_{2n+1}(ℤ) ⊗ ℤ_{l} = lim_{←} K_{2n+1}(ℤ, ℤ/l^{k}),

where the inverse limit on the right hand side is taken over the reduction of coefficients in K-groups with finite coefficients. In the beginning of the eigthies Soulé introduced the following construction cf. [So2]. Let ξ_{l^{k}} be a primitive root of unity of order l^{k}, e.g., ξ_{l^{k}} = exp \frac{2πi}{l^{k}}. Consider the number 1−ξ_{l^{k}} which is a unit in the ring R = ℤ[\frac{1}{l}, ξ_{l^{k}}]. There is the canonical Bott element β = β(ξ_{l^{k}}) ∈ K_{2}(R, ℤ/l^{k}) which depends only on the choice of ξ_{l^{k}}. Using products in the K-groups we can consider the nth power β^{n} ∈ K_{2n}(R, ℤ/l^{k}) of the Bott element and the product map

∗: K_{1}(R) ⊗ K_{2n}(R, ℤ/l^{k}) ⟶ K_{2n+1}(R, ℤ/l^{k}).

Since K_{1}(R) = R^{×} is the group of units, we can construct an element

(1 − ξ_{l^{k}}) ∗ β^{n} ∈ K_{2n+1}(R, ℤ/l^{k}).

Taking transfer map:

T r : K_{2n+1}(R, ℤ/l^{k}) → K_{2n+1}(ℤ, ℤ/l^{k})

down to K-theory of the integers gives

c_{l^{k}} = T r((1 − ξ_{l^{k}}) ∗ β^{n}) ∈ K_{2n+1}(ℤ, ℤ/l^{k}).

It can be checked using properties of the transfer that the elements c_{l^{k}} are compatible under the reductions of coefficients, hence they give a well defined element

(3.3)  c_{l} = lim_{←} c_{l^{k}} ∈ K_{2n+1}(ℤ) ⊗ ℤ_{l}

which is usually called the Soulé cyclotomic element. In this way we obtained one element c_{l} for any odd prime l. This simple construction turned out to be very powerful in applications. Note that to define c_{l} we have used products and transfers, which are encoded in the topology of the K-theory spectrum in the sense of algebraic topology. In the eighties there was yet another construction due to Beilinson which gave a nontrivial element in the rational K-group K_{2n+1}(ℤ) ⊗ ℚ. Recently it has been proven by Beilinson and Deligne [BD] (see also [HW] for a carefull check of details) that the element of Beilinson and the element of Soulé are compatible under the obvious map

K_{2n+1}(ℤ) ⊗ ℚ ⟶ K_{2n+1}(ℤ) ⊗ ℚ_{l},

for any odd l. Since K_{2n+1}(ℤ) is finitely generated, the compatibility implies that there exists a nontorsion element b_{2n+1} in K_{2n+1}(ℤ), which for any odd l maps to c_{l} by the map induced by the imbedding of ℤ into ℤ_{l}. Slightly abusing the terminology in what follows we will keep calling b_{2n+1} the Beilinson element. The number b from the assumption of the theorem is by definition the index of the subgroup of K_{2n+1}(ℤ) generated by b_{2n+1}.

Step 2. A map from K-theory down to cyclotomic units

We will denote by E the cyclotomic field ℚ(ξ_{l^{k}}) for k > 0. Note that E is the field of fractions of the ring O_{E}. For a finite set S of prime ideals of O_{E} by O_{E,S} we denote the ring of S-integers in E. By definition, the ring O_{E,S} consists of fractions from E with denominators divisible by prime ideals from S. The symbol O_{E,S}^{×} is the notation for the group of units of the ring. The map

(3.4)  α: K_{2n+1}(ℤ) ⟶ O_{E,S_{k}}^{×}/O_{E,S_{k}}^{×l^{k}} ⊗ ℤ/l^{k}(n)

which we are about to define takes values in twisted units of the cyclotomic field E. We use étale cohomology for conventional reasons but the reader should keep in mind that in what follows it can be replaced by the continuous cochains cohomology of Galois groups similar to the group G from Theorem 2.1. The map α is the composition of the following maps:

(3.5)  \begin{matrix}K_{2n+1}(ℤ) \\ α^{0} & ↓ \\ (K_{2n+1}(ℤ) ⊗ ℤ_{l})/l^{k} \\ α^{1} & ↓ \\ H^{1}(ℤ[\frac{1}{l}]; ℤ_{l}(n + 1))/l^{k} \\ α^{2} & ↓ \\ H^{1}(O_{E,S_{k}}; ℤ_{l}(n + 1))/l^{k} \\ α^{3} & ↓ \\ H^{1}(O_{E,S_{k}}; ℤ/l^{k}(n + 1)) \\ α^{4} & ↓ & ≅ \\ O_{E,S_{k}}^{×}/O_{E,S_{k}}^{×l^{k}} ⊗ ℤ/l^{k}(n) \end{matrix}

The map labeled α_{0} is induced by the obvious

K_{2n+1}(ℤ) → K_{2n+1}(ℤ) ⊗ ℤ_{l}

which sends x to x ⊗ 1. The map α_{1} is induced by the Dwyer-Friedlander map from K-theory to étale cohomology. Note that α_{1} is an isomorphism for prime numbers l such that K_{2n+1}(ℤ) = 0 has no l-torsion by our assumption on l. To define α_{2} we choose S_{k} to be any finite set of primes of E such that it contains primes over l and that the class group of the ring O_{E,S_{k}} vanishes. This can be done because the class group of every field is zero and Cl(E) = lim_{→} Cl(O_{E,S}), where the direct system is over all finite sets S of prime ideals of O_{E}. The arrow α_{2} is the injection which comes from the Hochschild-Serre spectral sequence of the monomorphism ℤ[\frac{1}{l}] → O_{E,S_{k}}. The map α_{3} is an imbedding which comes from the exact sequence in cohomology induced by the exact sequence

0 ⟶ ℤ_{l}(n+1) ⟶^{×l^{k}} ℤ_{l}(n+1) ⟶ ℤ/l^{k}(n+1) ⟶ 0.

Finally, the map α_{4} is the inverse of the isomorphism provided by tensoring the Kummer exact sequence

0 ⟶ O_{E,S_{k}}^{×}/O_{E,S_{k}}^{×l^{k}} ⟶ H^{1}(O_{E,S_{k}}; ℤ/l^{k}(1)) ⟶ Cl(O_{E,S_{k}}) ⟶ 0,

by ℤ/l^{k}(n), since Cl(O_{E,S_{k}}) = 0.

Lemma 3.6. The map

α: K_{2n+1}(ℤ) ⟶ O_{E,S_{k}}^{×}/O_{E,S_{k}}^{×l^{k}} ⊗ ℤ/l^{k}(n)

sends the Beilinson element b_{2n+1} onto u_{k}^{(n)} ⊗ ξ_{l^{k}}^{⊗n}, where u_{k}^{(n)} is the number

u_{k}^{(n)} = ∏_{1≤a<l^{k}, (a,l)=1} (1 − ξ_{l^{k}}^{a})^{a^{n}}.

Proof of Lemma 3.6 is a routine exercise. The key point is the well-known property of the transfer map: T r when composed with the map α_{2} gives the product over the elements of the Galois group Gal(E/ℚ). The exponents a in the formula for u_{k}^{(n)} come from the Galois action on the units in O_{E,S_{k}}, while the a^{n}’s indicate the Galois action on the module ℤ/l^{k}(n) = ℤ/l^{k}(1)^{⊗ n}. The numbers u_{k}^{(n)} provide the link between K_{∗}(ℤ) and the conjectures in number theory which we are going to discuss in the next section.

Step 3. Application of the theorem of Chebotarev We use the Chebotarev density theorem which enables us to count the density of the set M_{k}. We prepare the setup accordingly. The map α which we have defined by the diagram (3.5) appears as the left vertical arrow in the following commutative diagram.

(3.7)  \begin{matrix}K_{2n+1}(ℤ) & ⟶^{ϕ_{p}} & K_{2n+1}(𝔽_{p})_{l} = ℤ/l^{k} \\ ↓ & & ↓^{=} \\ (K_{2n+1}(ℤ) ⊗ ℤ_{l})/l^{k} & ⟶ & (K_{2n+1}(𝔽_{p}) ⊗ ℤ_{l})/l^{k} \\ ↓ & & ↓ \\ O_{E,S_{k}}^{×}/O_{E,S_{k}}^{×l^{k}} ⊗ ℤ/l^{k}(n) & ⟶ & κ_{w}^{×}/κ_{w}^{×l^{k}} ⊗ ℤ/l^{k}(n) \end{matrix}

The horizontal arrows in the diagram are induced by the reduction at p. Here w ∉ S_{k} is a prime of O_{E} which divides the principal ideal (p) and κ_{w} = O_{E,S_{k}}/w denotes its residue field. It is a finite field containing 𝔽_{p}. The lower horizontal arrow is the natural map induced by the projection O_{E,S_{k}} ⟶ κ_{w}. We will need a bit more arithmetic of the numbers u_{k}^{(n)}. Let L = E( \sqrt[l^{k}]{u_{k}}) be the field extension obtained by adding to E the l^{k}th root of u_{k}^{(n)}. The field is well defined because the l^{k}th roots of unity are in E. It is a cyclic extension of E with the Galois group G(L/E) = ℤ/l^{k}. We choose a prime ideal w̃ in O_{L} dividing w. The situation is shown at the picture below, where at the left hand side stands the tower of the field extensions. At the right we have indicated the chosen prime ideals such that w̃|w and w|(p).

L = E( \sqrt[l^{k}]{u_{k}})  w̃ E = ℚ(ξ_{l^{k}})  w ℚ  (p)

Inside the Galois group G(L/E) there exists the canonical automorphism F r_{w}, called the Frobenius automorphism, which generates the cyclic group Gal(κw̃/κ_{w}) ⊂ G(L/E). We have the following result.

Lemma 3.8. In this setting, the number of elements of the image of the reduction map ϕ_{p} equals the order of the Frobenius element F r_{w}.

Theorem of Chebotarev. Let L/K be a finite extension of number fields with Galois group G. Assume that the element σ ∈ G has c conjugates in G. Let P_{σ} denote the set of prime ideals of O_{K} which have a prime divisor in O_{L} whose Frobenius automorphism is σ. Then the set P_{σ} has the density which equals \frac{c}{#G}.

To finish the proof of Theorem 3.2 one uses the Chebotarev theorem for the extension L/ℚ. The Galois group G(L/ℚ) is the semidirect product of two cyclic groups: Gal(E/ℚ) = (ℤ/l^{k})^{×} and G(L/E) = ℤ/l^{k}. It is possible to compute the number of conjugates in G(L/ℚ) of the elements of a given order. Together with the Chebotarev theorem and Lemma 3.8 this implies the claim on the density of M_{k}. For more details we refer the interested reader to [G]. Proofs of the theorem of Chebotarev which can be found in textbooks on number theory, e.g., [Ne] usually rely on analytic methods. We recommend the survey paper [SL] which contains a sketch of the very first proof (purely algebraic) together with some interesting historical remarks about Chebotarev and his theorem.

4. Vandiver conjecture and K-theory

Let A denote the l-Sylow subgroup of the class group of the ring ℤ[ξ_{l}]. It has an action of the Galois group Δ := Gal(ℚ(ξ_{l})/ℚ) ≅ (ℤ/l)^{×}. We define the Teichmüller character

ω : Δ ⟶ (ℤ/l)^{×}

by σ(ξ_{l}) = ξ_{l}^{ω(σ)}, for σ ∈ Δ. Note that ω generates the character group

Hom(Δ, ℂ^{×}) ≅ ℤ/(l−1)

if we identify (ℤ/l)^{×} with the group of (l−1)th roots of unity μ_{l−1} ⊂ ℂ^{×}. Since Δ has order prime to l, we have the canonical decomposition cf. [Wa, section 6.3]

(4.1)  A = ⨁_{0≤i≤l−2} A^{[i]},

where

A^{[i]} = {a ∈ A : σ(a) = ω^{i}(σ)a for every σ ∈ Δ}

is the ith eigensubmodule of A. Note that A^{[i]} = e_{i}A, where

e_{i} = \frac{1}{l−1} ∑_{σ∈Δ} ω^{−i}(σ)σ

is the orthogonal idempotent of the ring ℤ_{l}[Δ]. It is well known that A^{[0]} = A^{[1]} = 0 cf. [W, Prop. 6.16].

Vandiver Conjecture. A^{[i]} = 0 for every even integer i such that 2 ≤ i ≤ l−3.

This statement was formulated by Vandiver about 70 years ago. Vandiver writes in [Va] that the conjecture had appeared to him much earlier, around 1912. It is also being attributed to Kummer, who considered the same statement on class numbers of cyclotomic fields in a letter to Kronecker dated on the 28th of December 1848, cf. [Ku, p. 84]. The Vandiver conjecture has been checked numerically for many primes. The most recent published account [BCEMS] shows that there is no counterexample for the conjecture for primes smaller than 12, 000, 000. Note that there exists an heuristic argument against the conjecture based on certain assumption on the divisibility properties of the numbers N_{n} by prime numbers [Wa, p. 158].

The numbers u_{k}^{(n)} which have been introduced in the previous section are of interest for number theorists because of their direct relation to the Vandiver conjecture.

Proposition 4.2. Let l be an odd prime and let n be an even interger such that 2 ≤ n ≤ l−3. Let k be a positive integer. The following statements are equivalent.

(1) The number

u_{k}^{(n)} = ∏_{(a,l)=1; 1≤a<l^{k}}(1 − ξ_{l^{k}}^{a})^{a^{n}}

can not be written as u_{k}^{(n)} = η^{l}, for any η ∈ ℤ[ξ_{l^{k}}].

(2) The eigensubmodule A^{[l−1−n]} is trivial.

Proof. We show how to reduce the proof to the similar statement for u_{1}^{(n)} which was obtained earlier by Thaine [Th]. To simplify notation a little bit, let u_{k}=u_{k}^{(n)}. For every k>1 the following three statements are equivalent

(1) u_{1} is an lth power in the group ℚ(ξ_{l})^{×}

(2) u_{1} is an lth power in the group ℚ(ξ_{l^{k}})^{×}

(3) u_{k} is an lth power in the group ℚ(ξ_{l^{k}})^{×}. In order to see this, observe that for every n and k we have the following equalities in the group ℚ(ξ_{l^{k}})^{×}/ℚ(ξ_{l^{k}})^{×l}.

u_{k} ≡ ∏_{1≤b<l} [  ∏_{a≡b mod l; 1≤a<l^{k}}(1 − ξ_{l^{k}}^{a})]^{b^{n}}

≡ ∏_{1≤b<l} [ ∏_{1≤j≤l^{k−1}}(1 − ξ_{l^{k}}^{b+lj})]^{b^{n}}

≡ ∏_{1≤b<l} [ ∏_{λ^{l^{k−1}} =1}(1 − λξ_{l^{k}}^{b})]^{b^{n}} (4.3)

≡ ∏_{1≤b<l} [(1 − (ξ_{l^{k}}^{b})^{l^{k−1}})]^{b^{n}}

≡ ∏_{1≤b<l}(1 − ξ_{l}^{b})^{b^{n}} = u_{1}

This shows that (2) and (3) are equivalent. Also, (1) clearly implies (2). To check that (2) implies (1), let us assume that u_{1} ∉ ℚ(ξ_{l})^{×l}. Consider the Kummer extension L = ℚ(ξ_{l},\sqrt[l]{u_{1}}). By assumption it has degree l over ℚ(ξ_{l}). It is not difficult to check that the Galois group G(L/ℚ) is not abelian. Since G(ℚ(ξ_{l^{k}})/ℚ) is obviously abelian, we see that L ⊄ ℚ(ξ_{l^{k}}), hence \sqrt[l]{u_{1}} ∉ ℚ(ξ_{l^{k}}). □

Using properties of the map α from (3.5) one can show the following result.

Theorem 4.4. [G, Th. 2.14] Assume that the l-torsion part of the group K_{2n+1}(ℤ) is trivial. Let l^{k_{0}} be the largest power of l which divides the index of the subgroup of K_{2n+1}(ℤ) generated by the Beilinson element. Let k be a natural number which is not smaller than k_{0}. If u_{k}^{(n)} = η^{l^{k_{1}}} , for some η ∈ ℤ[ξ_{l^{k}}], then k_{1} ≤ k_{0}.

Observe that the statement of Theorem 4.4 relies on topology, because in order to define the Soulé elements one needs products and transfers in K-groups, which come from the multiplicative structure of the K-theory ring spectrum. To the best of the author’s knowledge no bouds on the divisibility of the numbers u_{k}^{(n)} has been obtained by other methods.

Corollary 4.5. Fix a positive, even integer n. If the Quillen-Lichtenbaum conjecture is true, then the following two statements are equivalent.

(1) The group A^{[l−1−n]} = 0, for every odd prime l.

(2) The Beilinson element b_{2n+1} generates the group K_{2n+1}(ℤ) modulo 2-torsion.

The class groups of the cyclotomic field are also connected with the even dimensional K-groups, because of Theorem 2.1 and the following observation due to Kurihara, [K, Cor.1.5].

Proposition 4.6. Let l be an odd prime. If n > 0 is even, then A^{[l−1−n]}=0 iff H^{2}(G, ℤ_{l}(n+1))=0.

Note that the Galois twisting of the G-module ℤ_{l}(n+1) forced the unusual indexing of the eigensubmodules of A. Section 6 contains the proof of Proposition 4.6. Since K_{4}(ℤ)=0, and the group surjects onto H^{2}(G, ℤ_{l}(3)) by Theorem 2.1, we have

Corollary 4.7. [K, Cor.3.8] For any odd l the eigensubmodule A^{[l−3]} is zero.

This has not been proven by another method. A couple of years ago Soulé was able to extend the result of Kurihara cf. [So4].

Theorem 4.8. Assume that m > 1 is odd.^{1} If log l > m^{224m^{4}}, then A^{[l−m]} = 0.

The starting point of Soulé’s proof of Theorem 4.8 is as in [K]. By Theorem 2.1 we know that to control A^{[l−m]} it is enough to bound the torsion of the group K_{2m−2}(ℤ). Then Soulé uses the Hurewicz map and the stabilization theorem for the homology of the general linear group to get the map

K_{2m−2}(ℤ) ⟶ H_{2m−2}(SL_{N}(ℤ) , ℤ)

with N sufficiently large. It follows by a result of Arlettaz [A] that the kernel of the latter map has the exponent divisible only by primes smaller than m. Thus it is enough to bound the torsion of the homology of the SL_{N}(ℤ). The classical Voronoi “reduction theory” gives an explicit cell decomposition of the compactification of the locally symmetric space attached to the group SL_{N}(ℤ). With this in hand, Soulé implements the following brilliant observation of Gabber: there is an upper bound on the torsion in the homology of a finite CW-complex. The bound depends only on the data associated to the chain complex C_{∗}(X) of X, such as the number of cells of a fixed dimension and the number of faces of each cell. This yields the explicit bound on the torsion in K_{2m−2}(ℤ) and consequently the double exponential bound in Theorem 4.8. In the paper [So5, 5.5] Soulé sketches a way to prove that A^{[l−5]}=0, by bounding from below the value of the Borel regulator on K_{9}(ℤ).

As far as we are aware Theorem 4.4, Corollary 4.7 and Theorem 4.8 are the first results on the Vandiver conjecture of the general nature.

We have two additional reformulations of the Vandiver conjecture in terms of Ktheory.

1Put m = n+1 to return to our previous indexing.

Theorem 4.9. [BG1, section 5.2] The following statements are equivalent for a fixed odd prime l

(1) The Vandiver conjecture holds true for l.

(2) The groups of infinitely l-divisible elements D_{n+1}(ℚ)_{l} := ⋂_{k≥1} l^{k}K_{2n}(ℚ) of the K-groups of ℚ vanish for all n, even and positive.

(3) The étale K-theory group K_{4}^{et}(R) is trivial, where R=ℤ[\frac{1}{l}, ξ_{l}+ξ_{l}^{−1}] is the

ring obtained from ℤ by inverting l and adding the number ξ_{l}+ξ_{l}^{−1}=2 cos \frac{2π}{l} . If the Quillen-Lichtenbaum holds true, then D_{n+1}(ℚ)_{l}=K_{2n}(ℤ)_{l} in (2) and one can skip the superscript ét in the statement of (3). The l-torsion part of an abelian group C is denoted here by C_{l}.

Remark 4.10. The subgroup D_{n+1}(F)_{l} = ⋂_{k≥1} l^{k}K_{2n}(F) of infinitely l-divisible elements of K-groups of a number field F, for any n > 0 was introduced and investigated by G.Banaszak in [Ba1] and [Ba2]. Theorem 1.4 (1) implies that the group is contained in K_{2n}(O_{F}), hence it is finite. For k large enough the group is isomorphic to the cokernel of the boundary map ∂ in the localization sequence for the K-theory with finite coefficients:

K_{2n+1}(F, ℤ/l^{k}) ⟶^{∂} ⨁_{𝔭 |̸l} K_{2n}(κ_{𝔭}, ℤ/l^{k}) ⟶ D_{n+1}(F) ⟶ 0

(compare this sequence with the sequence (1.2) defining Cl(O_{F})). Because of the properties the group D_{n+1}(F) can be considered as an anologue in higher K-theory of O_{F} of the class group. For more on the group of infinitely divisible elements and its rich arithmetic the reader is refered to the papers: [Ba1], [Ba2], [BG1], [BG2] and [BGKZ]. Note that in the paper [BGKZ] the Quillen-Lichtenbaum conjecture was reformulated in terms of divisibility properties of elementary matrices.

If the Quillen-Lichtenbaum conjecture is proven, then the groups K_{2n+1}(ℤ) and K_{2n}(ℤ), with n>0 even, will be related like units and the class group of the cyclotomic field in a classical formula of Kummer. The formula of Kummer relates the index of the cyclotomic units in the group of units to the class number of the cyclotomic field ℚ(ξ_{l}) cf. [Wa, p. 145].

Proposition 4.11. Assume that the Quillen-Lichtenbaum conjecture holds true. Let n > 0 be an even integer and denote by ⟨b_{2n+1}⟩ the subgroup in K_{2n+1}(ℤ) generated by the Beilinson element b_{2n+1}. Then up to a power of 2 the following index equality

[K_{2n+1}(ℤ) : ⟨b_{2n+1}⟩] = #K_{2n}(ℤ)

holds true.

Proposition 4.11 is a consequence of the Main Conjecture cf. [BK, (6.6), p.385]. Observe that it shows that two approaches toward the Vandiver conjecture which were discussed in this section are in principle equivalent, if the Quillen-Lichtenbaum conjecture is true.

5. Iwasawa cyclicity conjecture

Theorem 5.1. [Herbrand-Ribet] Fix an odd prime number l. Let i be an odd integer such that 1 ≤ i ≤ l−2. Then the following two statements are equivalent.

(1) The group A^{[i]} is nontrivial.

(2) The prime l divides the number N_{l−i−1}.

Recall that in our notation N_{n} is the numerator of the divided Bernoulli number \frac{B_{n+1}}{n+1} . The implication (1) ⇒ (2) was known classically, [Wa, Thm. 6.17]. The converse and its proof due to Ribet is the cornerstone of the arithmetic of cyclotomic fields [Ri]. Nowadays this can be proven using the important techniques of Euler systems due to Kolyvagin and Rubin. The original proof of Ribet and its reinterpretation by Wiles [Wi] was a significant clue in the work of Mazur and Wiles on the Main Conjecture in [MW]. This in turn has direct consequences for the eigensubmodules A^{[i]}. For n as above we have by [Wa, Cor.5.15] \frac{B_{l−i}}{l−i} ≡ B_{1,ω^{−i}} mod l, where

B_{1,ω^{−i}} = \frac{1}{l} ∑_{a=1}^{l−1} aω^{−i}(σ_{a})

is a certain l-adic integer (the generalized Bernoulli number attached to the character ω^{−i}) and σ_{a} ∈ Δ is the automorphism maped by ω onto a ∈ ℤ/(l−1). The following result was proven by Wiles in [Wi].

Theorem 5.2. [Wiles] Let i be an odd integer such that 1 ≤ i ≤ l−2. Then A^{[i]} has exactly l^{k_{i}} elements, where l^{k_{i}} is the maximal power of l dividing B_{1,ω^{−i}}.

Iwasawa Cyclicity Conjecture. For every odd integer i such that 1 ≤ i ≤ l−2 the group A^{[i]} ≅ ℤ/l^{k_{i}}.

This conjecture appeared as an assumption of a theorem of Iwasawa cf. [I1, p.78]. The claim of the theorem concerned Galois module properties of class groups of towers of cyclotomic fields and was reformulated by Iwasawa in [I2] as a conjecture, later named the Main Conjecture in Iwasawa theory. The Main Conjecture for ℚ was proven by Mazur and Wiles in [MW] by considering reductions of modular curves. There exists more elementary proof of the Main Conjecture based on the method of Euler systems which is due to Kolyvagin and Rubin. We refer the interested reader to [C] and [Gr] for the introduction to the Iwasawa theory and to [Wa, Chapter 15] for the discussion of the Rubin’s proof of the Main Conjecture. It is well-known that the Vandiver conjecture implies the Iwasawa cyclicity conjecture [Wa, Cor. 10.15]. Contrary to the case of the Vandiver conjecture, there is no doubt among the experts that the cyclicity conjecture should hold true. When proven, it will simplify the arithmetic of cyclotomic fields considerably.

As far as the connection with Galois cohomology and K-theory is concerned we have the observation of Kurihara (for the proof see Section 6).

Proposition 5.3. [K, Cor. 1.5] Let l be an odd prime. If n > 0 is odd, then A^{[l−1−n]} is a cyclic group iff H^{2}(G, ℤ_{l}(n+1)) is a cyclic group.

Remark 5.4. In [BG1] new elements in K-groups of ℚ were constructed using certain Gauss sums introduced by Coates [C]. The elements have the Euler system property in the sense of Kolyvagin, which made them useful for computing the order of the subgroup D_{n+1}(ℚ) ⊂ K_{2n}(ℤ)_{l}, for n odd cf. [BG1, Th C]. In [BG2, Th. 2.4] the same elements were used to construct nonzero cohomology classes in the groups H^{2}(G, ℤ_{l}(n+1)), when n is odd. Assuming the Quillen-Lichtenbaum conjecture we obtain the following statement about the group K_{2n}(ℤ).

Theorem 5.5. Assume that l and n are as in Proposition 5.3 and that the Quillen-Lichtenbaum conjecture is true. Let k be so large that K_{2n}(ℤ)_{l} = K_{2n}(ℤ)[l^{k}]. Then every nonzero element x ∈ K_{2n}(ℤ)_{l} can be written in the form (note the resemblance to the Soulé’s element c_{l}):

x = B[T r(λ_{x} ∗ β^{n})],

where β = β(ξ_{l^{k}}) ∈ K_{2}(ℤ[ξ_{l^{k}}], ℤ/l^{k}) is the Bott element, T r denotes the transfer map in K-theory and

B: K_{2n+1}(ℤ, ℤ/l^{k}) ⟶ K_{2n}(ℤ)[l^{k}]

is the Bockstein map in K-theory with finite coefficients. Here λ_{x} is an element of K_{1}(ℚ(ξ_{l^{k}})), which depends on x. It was defined using Gauss sums cf. [BG2, p.11] .

Remark 5.6. Let n and l be odd. In the moment one of the most important problems concerning K_{∗}(ℤ) is to construct new nonzero elements - possibly generators, in the groups K_{2n}(ℤ)_{l}. In this respect Theorem 5.5 is not satisfactory enough because the Gauss sum λ_{x} and its arithmetical properties depend strongly on x. In the case of K_{2n+1}(ℤ)_{l}, (n and l odd) the nonzero elements came from algebraic topology, or more precisely from the image of the J-homomorphism cf. Remark 2.3. It is possible that in the case of K_{2n}(ℤ)_{l}, the expected nonzero elements will be constructed by a topological method. With the development of motivic cohomology and related theories (for example the algebraic cobordism theory of Voevodsky cf. [Vo2] and [LM]), such a construction may be accomplished in the future.

6. Proof of (4.6) and (5.3)

Proposition 6.1. Let l be an odd prime.

(1) If n > 0 is even, then A^{[l−1−n]}=0 iff H^{2}(G, ℤ_{l}(n+1))=0.

(2) If n > 0 is odd, then A^{[l−1−n]} is a cyclic group iff H^{2}(G, ℤ_{l}(n+1)) is a cyclic group.

Proof. Since the groups in question are finite l-groups, it is enough to show that for any l and n

(6.2)  A^{[l−1−n]}/l ≅ H^{2}(G, ℤ_{l}(n+1))/l.

Taking the long exact sequence in Galois cohomology associated to the sequence of Galois modules

0 ⟶ ℤ_{l}(n+1) ⟶^{×l} ℤ_{l}(n+1) ⟶ ℤ/l(n+1) ⟶ 0

we obtain that H^{2}(G, ℤ_{l}(n+1))/l = H^{2}(G, ℤ/l(n+1)), becuase H^{3}(G, ℤ_{l}(n+1))=0, by the cohomological dimension. In the remaining part of the proof it is more convenient to replace the Galois cohomology by the isomorphic etale cohomology groups. In particular, we are going to prove that

(6.3)  A^{[l−1−n]}/l ≅ H^{2}(ℤ[\frac{1}{l}]; ℤ/l(n + 1)).

Let R be the ring ℤ[\frac{1}{l}, ξ_{l}] and let F denote the cyclotomic field ℚ(ξ_{l}). Consider the exact sequence in etale cohomology

0 ⟶ H^{1}(R, 𝔾_{m})/l ⟶ H^{2}(R, μ_{l}) ⟶ H^{2}(R, 𝔾_{m})[l] ⟶ 0

which comes from the exact sequence of group schemes over Spec R

1 ⟶ μ_{l} ⟶ 𝔾_{m} ⟶^{×l} 𝔾_{m} ⟶ 1.

Note that μ_{l}=ℤ/l(1). By the basics of etale cohomology cf. [Mi] we know that H^{1}(R, 𝔾_{m})=P ic(R) is the class group of the ring R and H^{2}(R, 𝔾_{m})=Br(R) is its Brauer group. It follows by the class field theory that Br(R)[l]=0, the reason being that there is a single prime ideal in R above l (recall that in R we have the equality of ideals (l)=(1−ξ_{l})^{l−1}). Hence we get

H^{2}(R, μ_{l}) = H^{1}(R, 𝔾_{m})/l = A/l = P ic(R)/l.

Tensoring with ℤ/l(n) and taking invariants under the action of Δ gives

(A ⊗ ℤ/l(n))^{Δ} ≅ H^{2}(R, ℤ/l(n+1))^{Δ}.

Since Δ has order prime to l, it follows by the Hochschild-Serre spectral sequence that the group on the right hand side is isomorphic to H^{2}(ℤ[\frac{1}{l}]; ℤ/l(n+1)). The group on the left hand side equals A^{[l−1−n]}/l, what can be checked by a simple computation. □

References

[A] D. Arlettaz, The Hurewicz homomorphism in algebraic K-theory, Jour. of Pure and Appl. Algebra 71 (1991), 1-12.

[ABG] D. Arlettaz, G. Banaszak, W. Gajda, On 2-adic cyclotomic elements in K-theory and étale cohomology of the ring of integers, the Journal of Number Theory 82 (2000), 225-255.

[Ba1] G. Banaszak, Algebraic K-theory of number fields and rings of integers and the Stickelberger ideal, Annals of Math 135 (1992), 325-360.

[Ba2] G. Banaszak, Generalization of the Moore exact sequence and the wild kernel for the higher K-groups, Compositio math. 86, No.3 (1993), 281-305.

[BCEMS] J. Buhler, R. Crandall, R. Ernvall, T. Metsälankylä, M.A. Shokrollahi, Irregular primes and cyclotomic invariants to twelve million, J. Symbolic Computation 11 (1999), 1-8.

[BD] A. Beilinson, P. Deligne, Motivic polylogarithm and Zagier conjecture, preprint (1992).

[Be] A. Beilinson, Higher regulators and values of L-functions, Jour. Soviet Math. 30 (1985), 2036-2070.

[BG1] G. Banaszak, W. Gajda, Euler systems for higher K-theory of number fields, Jour. of Number Theory 58, No. 2 (1996), 213-252.

[BG2] G. Banaszak, W. Gajda, On the arithmetic of cyclotomic fields and the K-theory of ℚ., Contemp. Math. 199 (1996), 7-18.

[BGKZ] G. Banaszak, W. Gajda, P. Krasoń, P. Zelewski, A note on the Quillen-Lichtenbaum conjecture and the arithmetic of square rings, K-theory 16 (1999), 229-243.

[BK] S. Bloch, K. Kato, L-functions and Tamagawa numbers of motives, in P. Cartier et al. (eds.) “The Grothendieck Festschrift” I (1990), Birkhaüser, 333-400.

[Bo] A. Borel, Cohomologie de SL_{n} et values de fonctiones zeta, Ann. Acad. Scuola Normale Superiore 7 (1974), 613-636.

[Br] W. Browder, Algebraic K-theory with coefficients. In: Geometric applications of homotopy theory I, Evanston 1977, Lecture Notes in Mathematics 658 (1974), 40-84.

[C] J. Coates, p-adic L-fuctions and Iwasawa theory In: Algebraic Number Fields, Durham 1975 (1977), Academic Press, 269-353.

[DF] W. Dwyer, E. Friedlander, Algebraic and étale K-theory, Trans. Amer. Math. Soc. 292 (1985), 247-280.

[E-VGS] P. Elbaz-Vincent, H. Gangl, C. Soulé, Quelques calculus de la cohomologie de GL_{N} (ℤ)

et de la K-théorie de ℤ, C. R. Math. Acad. Sci. Paris 335 (2002), 321-324.

[G] W. Gajda, On cyclotomic numbers and the reduction map for the K-theory of the integers, K-theory 23 (2001), 323-343.

[GL] T. Geisser, M. Levine, The Bloch-Kato conjecture and a theorem of Suslin-Voevodsky, J. Reine Angew. Math. 530 (2001), 55-103.

[Gr] R. Greenberg, Iwasawa theory-past and present, Advanced Studies in Pure Mathematics 9 (2001), 407-464.

[HW] A. Huber, J. Wildeshaus, Classical motivic polylogarithm according to Beilinson and Deligne, Doc. Math. 3 (1998), 27-133, ibidem 297-299.

[I1] K. Iwasawa, On some modules in the theory of cyclotomic fields, J. Math. Soc. Japan 16 (1964), 42-82.

[I2] K. Iwasawa, Analogies between number fields and function fileds. In: Some Recent Advances in Basic Sciences 2 (1969), Yeshiwa University, 203-208.

[K] M. Kurihara, Some remarks on conjectures about cyclotomic fields and K-groups of ℤ, Compositio Mathematica 81 (1992), 223-236.

[KFN1] M. Kolster, V. Fleckinger, T. Nguyen Quang Do, Twisted S-units, p-adic class number formulas and the Lichtenbaum conjectures, Duke Math. J. (1996), 679-717.

[KFN2] M. Kolster, V. Fleckinger, T. Nguyen Quang Do, Correction to “Twisted S-units, padic class number formulas and the Lichtenbaum conjectures”, Duke Math. J. (1997), 641-643.

[Ku] E.E. Kummer, Collected Papers vol. I, Springer, 1975.

[La] S. Lang, Algebra, Springer, 2002.

[Le1] M. Levine, Relative Milnor K-theory, K-theory 6 (1992), 113-175.

[Le2] M. Levine, Correction to: “Relative Milnor K-theory”, K-theory 9 (1995), 503-505.

[Li] S. Lichtenbaum, Values of zeta function, étale cohomology, and algebraic K-theory. In: Algebraic K-theory II, Seattle 1972, Lecture Notes in Math. 342 (1973), 489-501.

[LM] M. Levine, F. Morel, Algebraic cobordism I, preprint (2001).

[Lo] J-L. Loday, K-théorie algébraique et repré sentations de groupes, A.. Sci. Ecol. Nor. Sup. (4) 9 (1976), 309-377.

[LS1] R. Lee, R.H. Szczarba, The group K_{3}(ℤ) is cyclic of order forty-eight, Annals of Math. 104 (1976), 31-60.

[LS2] R. Lee, R.H. Szczarba, On the torsion in K_{4}(ℤ) and K_{5}(ℤ) with an Addendum by C. Soulé, Duke Math. J. 45 No.1 (1978), 101-132.

[M] J. Milnor, Introduction to algebraic K-theory, Princeton University Press, 1971.

[MW] B. Mazur, A. Wiles, Class fields of abelian extensions of |Q, Invent. math. 76 (1984), 179-330.

[Mi] J.S.Milne, Arithmetic Duality Theorems, Perspectives in Math., vol.1, Academic Press, 1986.

[N] J. Neisendorfer, Primary homotopy theory, Mem. Amer. Math. Soc. 232 (1980).

[Ne] J. Neukirch, Algebraische Zahlentheorie, Springer Verlag, 1997.

[Q1] D. Quillen, Higehr Algebraic K-theory I. In: Higher K-Theories, Seattle 1972, Lecture Notes in Math. 341 (1973), Springer, 85-147.

[Q2] D. Quillen, On the cohomology and K-theory of the general linear group over finite field, Ann. of Math. 96 (1972), 552-586.

[Q3] D. Quillen, Finite generation of the groups K_{i} of rings of algebraic integers In: Higher K-Theories, Seattle 1972, Lecture Notes in Math. 341 (1973), Springer, 179-198.

[Q4] D. Quillen, Letter from Quillen to Milnor on Im(π_{i}(O) → π_{i}^{s} → K_{i}(ℤ)). In: Algebraic K-theory, Evanston 1976, Lecture Notes in Math. 551 (1976), Springer, 182-188.

[R] J. Rosenberg, Grad. Texts Math. 147 (1994), Springer.

[Ri] K. Ribet, A modular construction of unramified p-extensions of ℚ(μ_{p}), Invent. math. 34 (1976), 151-162.

[Ro] J. Rognes, K_{4}(ℤ) is a trivial group, Topology 39 (2000), 267-281.

[RW] J. Rognes, C. Weibel, Two-primary algebaric K-theory of integers in number fields, J. Amer. Math. Soc. 13 (2000), 1-54.

[Sch] P. Schneider, Über gewisse Galoiskohomologiegruppen, Math. Z. 168 (1979), 181-205.

[SL] P. Stevenhagen, H.W. Lenstra, Chebotarev and his density theorem, Math. Intelligencer 18 (1996), 26-37.

[So1] C. Soulé, K-theorié des anneaux d’entries de corps de nombres et cohomologie étale, Invent. math. 55 (1979), 251-295.

[So2] C. Soulé, Éléments cyclotomiques en K-théorie, Astérisque 148-149 (1987), 225-257.

[So3] C. Soulé, On higher p-adic regulators, Lecture Notes in math. 854 (1981), 372-401.

[So4] C. Soulé, Perfect forms and the Vandiver conjecture, J. Reine Angew. Math 517 (1999), 209-221.

[So5] C. Soulé, A bound for the torsion in the K-theory of algebraic integers, preprint (November 2002).

[So6] C. Soulé, On the 3-torsion in K_{4}(ℤ), Topology 39 (2000), 259-265.

[SV] A. Suslin, V. Voevodsky, Bloch-Kato conjecture and motivic cohomology with finite coefficients. In: The arithmetic of algebraic cycles, Banff, 1998 (2000), NATO Sci. ser. C. Math. Phys. Sci. vol. 548, Kluwer Verlag, 117-189.

[Th] F. Thaine, Polynomials generalizing binomial coefficients and their application to the study of Fermat’s Last Theorem, the Jour. of Number Theory 15 (1982), 304-317.

[Va] H.S. Vandiver, Fermat’s Last Theorem and the second factor in the cyclotomic numbers, Bull. Amer. Math. Soc. 40 (1934), 118-126.

[Vo1] V. Voevodsky, The Milnor Conjecture, preprint (1996).

[Vo2] V. Voevodsky, 𝔸^{1}-homotopy theory, Doc. Math. J. DMV (Extra volume ICM 1998) (1998), 579-604.

[Wa] L. Washington, Introduction to cyclotomic fields, Springer Verlag, 1997.

[Wal] F. Waldhausen, Algebraic K-theory of generalized free products, Annals of Math. 108 (1978), 135-256.

[We1] C. Weibel, A survey of products in algebraic K-theory. In: Algebraic K-theory, Evanston, 1980, Lecture Notes in Mth. 854 (1981), 494-517.

[We2] C. Weibel, The 2-torsion in the K-theory of the integers, C. R. Acad. Sci. Paris Ser. I 324 (1997), 615-620.

[Wi] A. Wiles, Modular curves and the class group of ℚ(μ_{p}), Invent. math. 58 (1980), 1-35.

Math. Dept., The Adam Mickiewicz University, Umultowska 87, 61614 Poznań, POLAND

   E-mail address: GAJDA@math.amu.edu.pl

Current address: Max-Planck-Institut für Mathematik, Bonn, GERMANY

   E-mail address: GAJDA@mpim-bonn.mpg.de